Need help thanks.
Homework: Chapter 1 Suggested Homework Save Score: 0 of 1 pt 170120 (16complete) ▼ Hw Score: 78.33%, 15.67 of 20 pts Problem 1.14 Question Help * Chari s Lackey operates a bakery in Idaho Fal lito. Because of its excellert produ a de celent lo ation, de nr d has increased by 35% in the last yea. O fr to many o atom a st have not been able to purchase the bread of their choice. Because of the size of the store, no new ovens can be added. At a staff meeting. one employee suggested ways to load the ovens differently so hat more loaves of bread can be baked at one time. This new process will require that the ovens be loaded by hand, requiring addtional manpower. This is the only production change that will be made in order to meet the increased demand. The bakery curently makes 1,600 loaves per month Employees are paid $8 per hour, in addition to the labor cost, Charles also has a constant utility cost per month of $700 and a per loaf ingredient cost of $0.40 Current multifactor productivity for 640 work hours per mort. □ loavesrdollar (round yourresponso to moo decinal places.
Answer #1
  • Utility cost = $700 per month

The number of loaves prepared = 1600 per month

  • Ingredient cost = $0.40 per loaf =1600 *0.4 = $640

Labour cost = $8 per hour

  • Cost for 640 work hours per month = $5120

Total Cost = Labour cost + Ingredient cost + utility cost = $5120 + $640 + $700 = $6460

Multifactor productivity in loaves per dollar = 1600 / 6460 = 0.2476 loaves / USD

MFP or TFP measures economic performance comparing the output (the amount of services & goods produced) to the combined inputs including capital, labour, materials, energy, and, purchased services used to produce the services & goods.